OpenStudy (anonymous):

Which of the following points lies on the circle whose center is at the origin and whose radius is 13? a(-5, 12) b(squareroot13,squareroot 13 ) c(6, -7)

OpenStudy (anonymous):

Not sure why you created a duplicate here.. Did you try the distance formula?

OpenStudy (anonymous):

i dont know.. i havent ever done this kind of question....

OpenStudy (anonymous):

What is the distance from \((0,0)\) to \((-5,12)\) ? What is the distance from \((0,0)\) to \((\sqrt{13},\sqrt{13})\) ? What is the distance from \((0,0)\) to \((6,-7)\) ? The distance formula between two points is: \[Distance = \sqrt{(x_2-x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

Where one of the points is \((x_1,y_1)\) and the other point is \((x_2,y_2)\)

OpenStudy (anonymous):

soo its the last one? (:

OpenStudy (anonymous):

Nope.. \[\sqrt{6^2 + (-7)^2} = \sqrt{36 + 49} = \sqrt{85} \ne 13\]